OpenStudy (anonymous):

what is the range of this function f(x) = x2, with domain [1, ∞)

jimthompson5910 (jim_thompson5910):

The smallest value you can plug in is x = 1 f(x) = x^2 f(1) = 1^2 f(1) = 1 So the smallest value in the range is y = 1 This means that the range is also [1, ∞)